# Imagination Policy: Using Generative Point Cloud Models for Learning Manipulation Policies

### v1

@@ Abstract
Humans can imagine goal states during planning and perform actions to match those goals. In this work, we propose Imagination Policy, a novel multi-task key-frame policy network for solving high-precision pick and place tasks. Instead of learning actions directly, Imagination Policy generates point clouds to imagine desired states which are then translated to actions using rigid action estimation. This transforms action inference into a local generative task. We leverage pick and place symmetries underlying the tasks in the generation process and achieve extremely high sample efficiency and generalizability to unseen configurations. Finally, we demonstrate state-of-the-art performance across various tasks on the RLbench benchmark compared with several strong baselines.Project website: https://haojhuang.github.io/imagine_page
@@ 1 Introduction
Humans can look at a scene and imagine how it would look with the objects in it rearranged. For example, given a flower and a bottle on the table, we can imagine the flower placed in the bottle. Using this mental picture, we can then manipulate the objects to match the imagined scene. However, most robotic policy learning algorithms [1, 2, 3, 4, 5, 6] directly map observations to actions, which does not highlight the importance of local geometric information. This limits transferability between tasks and different robots, requiring more demonstrations and hindering generalization to novel objects and scenes.
@@
Inspired by how humans solve tasks, we propose Imagination Policy which takes two point clouds as input and generates a new point cloud combining the inputs into a desirable configuration using a conditional point flow model, as shown in Figure 1. Given the generated point cloud, we use point cloud registration methods to match the observed input point clouds with the “imagined” scene. This gives rigid transformations which can be used to command a robot arm to manipulate the objects. Imagination Policy consists of two generative processes, each of which uses the above method. As shown in Figure 1a, the pick generator generates the points of the object positioned relative to the gripper point cloud. The place generator generates a pair of objects rearranged together as shown in Figure 1b. Note that although the full method produces rigid transformations, the generative point cloud model is not constrained to produce rigid transforms of the input point clouds; it is only trained to do so. Compared to directly generating actions, this adds many degrees of freedom to the generative process which aids optimization and sensitivity to geometric interactions.
Learning a single robot policy to solve various manipulation tasks is a fundamental goal of robot learning. However, it is a challenge to learn from multimodal robot data across different tasks. Recently, transformer-based manipulation policies [1, 2, 3] have shown impressive multi-task capabilities. In the meantime, diffusion-based policies [7, 4, 3, 8] have demonstrated compelling results in processing multimodal robot data. However, most of these methods require hundreds of expert demonstrations and fail on tasks requiring high precision. For example, methods like PerAct [1], RVT [2] and Diffuser Actor [3] struggle to solve RLBench tasks [9] like Plug-Charger or Insert-Knife. Instead, Imagination Policy amortizes the action prediction by estimating the drift force of each point and can be used to solve high-precision tasks with few demonstrations.
@@
The geometric nature of our method also allows us to incorporate task symmetries as constraints, further improving learning efficiency. Pick-place tasks are symmetric with respect to the group of 3D rotations $\mathrm{SO}(3)$. The outcome of a grasp attempt is invariant to rotations and translations which move both the grasped object and the gripper by the same amount. Place interactions have similar symmetries involving separate motions of the picked object and the placed object. Consider a task involving placing an object in a slot. Independent rotations of the object ($g_{1}$) and the slot ($g_{2}$) result in a change ($a^{\prime}=g_{2}ag_{1}^{-1}$) in the requisite action needed to perform the insertion. This is referred to as bi-equivariance. Previous works [10, 11, 12, 13] have explored bi-equivariant symmetries in manipulation policy learning and achieved dramatic improvements in sample efficiency. However, these are limited to single-task pick-and-place settings. As a result, they cannot be directly applied to Plug-Charger and Insert-Knife tasks without a predefined preplace action. In contrast, Imagination Policy achieves bi-equivariance in a multi-task setting with key-frame control. As shown in Figure 1a, rotations of the flower will result in the same pick pose. Similarly, seperate rotations of the flower and the mug in Figure 1b will not change the generated point cloud.
Our contributions are as follows: 1) We propose using generative point cloud models and rigid action estimation for learning key-frame manipulation policies; 2). We propose Imagination Policy, a multi-task manipulation policy network, which includes a pick generation module and a place generation module and realizes bi-equivariance in the key-frame setting; 3). We demonstrate state-of-the-art performance on various experiments against several strong baselines.
@@ 2 Related Work
Point Cloud Generation. Previous works have explored point cloud generation using VAEs [14, 15] and GANs [16, 17]. Recently, score-based denoising models and normalizing flows [18, 19, 20, 21, 22, 23] have demonstrated the power and flexibility to generate high-quality point clouds. For example, Zhou et al. [20] proposed a probabilistic diffusion approach for unconditional point cloud generation. [21, 22] formulated conditional point cloud diffusion. PSF [24] achieved fast point cloud generation with rectified flow [23]. Ours, however, generates pick and place point clouds conditioned on the observation that can be used to estimate a rigid action to command the robot arm.
@@
Point Clouds in 3D Pick-and-place Manipulation. Point clouds provide a flexible, geometric representation to encode object shapes and poses. In terms of pick-and-place manipulation, Simeonov et al. [25] used $\mathrm{SE}(3)$-invariant point features to encode descriptor fields enabling sample efficient policy learning. Simeonov et al. [4] extended Diffusion Policy [7] to work with point cloud observations and to learn multimodal actions. Pan et al. [26] propose TAX-POSE which is closely related to our method. Pan et al. [26] used two segmented point clouds as input and directly output a new point cloud using a weighted summary of target points together with residual predictions. Concurrently, Eisner et al. [27] adopted TAX-POSE [26] with relative distance inferred by a kernel method. However, they were designed to output the new point cloud in one step directly without any penalty on the generated results. Ours, instead, uses generative models to predict the movement of each point iteratively with a velocity model. Moreover, they are limited to single-task training, only work in one-step pick-and-place settings, and thus cannot be applied to complex tasks without predefined prior actions. Recently, Shridhar et al. [1], Goyal et al. [2], and Ke et al. [3] showed impressive multi-task capabilities with transformer-based architectures. However, these methods require hundreds of expert demonstrations and cannot successfully learn high-precision tasks. In contrast, our method leverages bi-equivariant symmetry and amortizes the action prediction across multiple tasks. As a result, it can solve high-precision tasks with few demonstrations.
@@
Symmetry in Robot Learning. Robotic tasks defined in 3D Euclidean space are invariant to translations, rotations, and reflections which redefine the coordinate frame but do not otherwise alter the task. Recent advancements in equivariant modeling [28, 29, 30, 31] provide a convenient tool to encode symmetries in robotics. Zhu et al. [32] and Huang et al. [33] utilized equivariant models to enforce pick symmetries for grasp learning. Other works [25, 10, 34, 11, 12, 13] leverage symmetries in pick and place and achieve high sample efficiency. However, they are limited to single-task pick-and-place equivariance. As a result, they cannot be directly applied to the Plug-Charger and Insert-Knife tasks without a pre-place action. Our proposed method, however, can achieve bi-equivariance in the key-frame and multi-task setting. In addition, we realize the equivariant action inference using an invariant point cloud generating process, which is different from previous methods.
@@ 3.1 Pair Generation for Place
The generator input of $p^{k}$ is constructed as $(x_{t}^{k},f^{k},f_{\ell},t)$. We then optimize $\theta$ with respect to the loss function defined in Equation 2.
Estimating the Action. Given two sets of points $P_{a}\cup P_{b}=(p_{a}^{1},p_{a}^{2},\cdots,p_{a}^{n})\cup(p_{b}^{1},p_{b}^{2},\cdots,p_{b}^{m})$ and their corresponding target positions $P_{ab}=(T_{a}p_{a}^{1},T_{a}p_{a}^{2},\cdots,T_{a}p_{a}^{n})\cup(T_{b}p_{b}^{1},T_{b}p_{b}^{2},\cdots,T_{b}p_{b}^{m})$, we can recover the rigid transformations $T_{a}$ and $T_{b}$ using SVD [35]. However, the output $\hat{P}_{a}\cup\hat{P}_{b}$ of the generator $f_{\mathrm{gen}}$ is not constrained to be given by rigid transforms of the original two point clouds. Each point may move independently by transformation $T^{i}_{\alpha}$ such that $\hat{P}_{a}\cup\hat{P}_{b}=(T^{1}_{a}p_{a}^{1},T^{2}_{a}p_{a}^{2},\cdots,T_{a}^{n}p_{a}^{n})\cup(T^{1}_{b}p_{b}^{1},T^{2}_{b}p_{b}^{2},\cdots,T^{m}_{b}p_{b}^{m})$. We can still use SVD to estimate the best fitting $\hat{T}_{a}$ between $(P_{a},\hat{P}_{a})$ as well as $\hat{T}_{b}$ between $(P_{b},\hat{P}_{b})$. Assuming $P_{b}$ represents the object to rearrangement and $P_{a}$ represents the placement, as shown in Figure 2, the pre-place or place action can be calculated as $a_{\mathrm{place}}=(\hat{T}_{a})^{-1}\hat{T}_{b}$.
Figure 2: Architecture of Imagination Policy. (a). Encoding the observed point features. (b). Conditional pair generation of the place scene from random Gaussian noise. (c). Estimating the rigid transformation from the observed point cloud to the generation using correspondence.
Realizing the Bi-equivariance. As noted in prior work [10, 11, 12], place actions that transform an object B with respect to another object A are bi-equivariant. That is, independent transformations of object B with $g_{b}\in\mathrm{SE}(3)$ and object A with $g_{a}\in\mathrm{SE}(3)$ result in a change ($a^{\prime}_{\mathrm{place}}=g_{a}a_{\mathrm{place}}g_{b}^{-1}$) to complete the rearrangement at the new configuration. Leveraging bi-equivariant symmetries can generalize the learned place knowledge to different configurations and improve sample efficiency.
@@
Assuming Gaussian noise $X_{0}$ is rotation-invariant, if the encoded point feature $F_{a}$ and $F_{b}$ are rotation-invariant, then $f_{\mathrm{place}}$ is bi-equivariant
@@
With the same generated points, the estimated transformation from rotated observation $g_{a}P_{a}={(g_{a}p_{a}^{1},g_{a}p_{a}^{2},\cdots,g_{a}p_{a}^{n})}$ to $\hat{P}_{a}$ is $\hat{T}_{a}g_{a}^{-1}$. Similarly, the estimated transformation from $g_{b}P_{b}$ to $\hat{P}_{b}$ is $\hat{T}_{b}g_{b}^{-1}$. Then, the new place action $a^{\prime}_{\mathrm{place}}$ can be calculated as $a^{\prime}_{\mathrm{place}}=(\hat{T}_{a}g_{a}^{-1})^{-1}\hat{T}_{b}g_{b}^{-1}=g_{a}\hat{T}_{a}^{-1}\hat{T}_{b}g_{b}^{-1}=g_{a}a_{\mathrm{place}}g_{b}^{-1}$, which satisfies bi-equivariance .∎
@@ 3.2 Single Generation for Pick
Assuming Gaussian noise $X^{P_{a}}_{0}$ is rotation-invariant, $f_{\mathrm{pick}}$ is rotation-equivariant if the encoded point feature $F_{a}$ and $F_{b}$ are rotation-invariant: $f_{\mathrm{pick}}(P_{a},g_{b}\cdot P_{b})=g_{b}f_{\mathrm{pick}}(P_{a},P_{b})$.
Specifically, if there is a rotation $g_{b}$ acting on $P_{b}$, the generated points $\hat{P}_{b}$ are the same as those without rotation. The estimated transformation from $g_{b}P_{b}$ to $\hat{P}_{b}$ is $\hat{T}_{b}g_{b}^{-1}$ and the new pick action can be calculated as $a^{\prime}_{\mathrm{pick}}=(\hat{T}_{b}g_{b}^{-1})^{-1}=g_{b}\hat{T}_{b}^{-1}$, which realizes the equivariance.
@@ 4 Experiments
Model Architecture Details. The generation models of $f_{\mathrm{pick}}$ and $f_{\mathrm{place}}$ share the same architecture. Each has two point cloud encoders and a generation network. We select PVCNN [37] as the backbone of our point encoders, which output a 64-dimension feature for each point. We use pre-trained CLIP-ViT32 [36] as our language encoder and project the language embedding to a 32-dimension vector with a linear layer. The time step t is encoded as a 32-dimension positional embedding. We also encode a binary mask that indicates if the point belongs to $P_{a}$ or $P_{b}$ as a 32-dimension positional embedding. As a result, the generator input of a point is a 163-dimension vector. We adopt PSF [24] as our generator backbone. Both $f_{\mathrm{pick}}$ and $f_{\mathrm{place}}$ are trained end-to-end with the MSE loss defined in Equation 2 and Equation 5. We use the Adam optimizer with an initial learning rate of $10^{-4}$. Training takes 7 hours to converge with $200$k training steps on a single RTX-4090 graphic card. During inference, we randomly sample $X_{0}$ from a Gaussian distribution and integrate over $v_{\theta}$ with 1000 steps to generate $P_{ab}$ and calculate the action. Generating one batch takes $20$ seconds.
All point clouds $P_{a}$ and $P_{b}$ in our experiments are captured by RGB-D cameras instead of directly sampling from the ground truth mesh. We first center the point cloud and then downsample by selecting at most one point in each cell of a 4mm voxel grid. We further randomly subsample or duplicate to get $2048$ points for $P_{a}$ and $P_{b}$. To get rotation-invariant generation, we apply extensive $\mathrm{SO}(3)$ data augmentation to $P_{a}$ and $P_{b}$ during training, i.e., an $\mathrm{SO}(3)$ rotation is sampled uniformly at each training step. It enforces $f_{\mathrm{gen}}(g_{a}P_{a},g_{b}P_{b})=f_{\mathrm{gen}}(P_{a},P_{b})$, which leads to the desired symmetry properties from Proposition 1. We found the results slightly outperform the equivariant point encoder of Vector Neurons [29], as shown in Table 2.
@@
Baselines. Our method is compared against three strong baselines: PerAct  [1] is the state-of-the-art multi-task behavior cloning agent that tokenizes the voxel grids together with a language description of the task and learns a language-conditioned policy with Perceiver Transformer [38]. RVT  [2] projects the 3D observation onto five orthographic images and uses the dense feature map of each image to generate 3D actions. 3D Diffuser Actor [3] is a variation of Diffusion Policy [7] that denoises noisy actions conditioned on point cloud features. Comparison with this baseline tests the importance of point cloud generation since this baseline generates actions directly.
Settings. All methods are trained as multi-task models as in their original settings. There are four cameras (front, right shoulder, left shoulder, hand) pointing toward the workspace. For our method, we formulate the action sequence as (pick, preplace, place), as shown in Figure 4. Specifically, our method generates the pick action with $f_{\mathrm{pick}}$, and the preplace and place action with $f_{\mathrm{place}}$ simultaneously. We use the ground truth mask to segment $P_{a}$ and $P_{b}$, as shown in Figure 4a.
@@
Results. We report the results of all methods in Table 1. Several conclusions can be drawn from Table 1: 1) Imagination Policy significantly outperforms all baselines trained with 10 demos on all the tasks except Put-Plate. It can also achieve over $90\%$ success rates in Phone-on-Base and Stack-Wine. 2) For tasks with a high-precision requirement, e.g., Plug-Charger, Insert-Knife and Put-Roll, Imagination Policy has a relatively high success rate while all the baselines fail to learn a good policy. 3) Imagination Policy achieves better sample efficiency and demonstrates few-shot learning performance. With one or five demonstrations, it sometimes outperforms the baselines trained with 10 demonstrations, e.g., Imagination Policy achieves a $97.2\%$ success rate on Stack-Wine trained with 5 demos while the best baseline can only achieve ${26.67\%}$. We believe this sample efficiency is due to the ability of equivariance to exploit the symmetry inherent in the generation. In the end, our method underperforms one baseline in Put-Plate. We hypothesize that the object in this task is symmetric and is hard to encode as distinguishable point features, which might result in wrong correspondences when estimating the rigid transformations.
@@
4.2 Ablation Study
Multimodal Pick-part Dataset. To quantitatively measure the point cloud generation results and the equivariance of Imagination Policy, we create a small pick-part dataset using four YCB objects [39] (banana, mug, spoon and fork). We load each object in the Pybullet simulator [40] and use three cameras to get the RGB-D images to extract the point cloud. Each object is assigned with two different expert grasps with corresponding language instructions, e.g., “grasp the mug by the handle”, “grasp the mug by its body”, as shown in Figure 6.
Training and Metrics. We trained a single pick generation model to generate all the objects conditioned on the canonicalized gripper points and language descriptions. We also train a place generation model to generate both the gripper point cloud and the object point cloud. To evaluate the pick generation results, we randomly rotate and randomly downsample the object point cloud ($P_{b}$) to make a starting pose unseen during training. We calculate the translation error and rotation error between the estimated grasp pose and the ground truth grasp pose. Note that rotating $P_{b}$ results in a change of the ground truth pick pose. To evaluate the place generation results, we randomly rotate and downsample the gripper ($P_{a}$) as well as the object ($P_{b}$) to make a scene unseen during training and calculate the translation error and rotation error between estimated transformation $\hat{T}_{a}^{-1}\hat{T}_{b}$ and the ground truth pose. Note that rotating either $P_{a}$ or $P_{b}$ changes the relative ground truth transformation. We report the minimum, mean and maximum error over 100 runs in Table 2. We also show visualizations of the generated point clouds in orange in Figure 6.
Results. Table 2 includes 6 variations of our proposed methods. Several findings can be concluded from Table 2: (1) As shown in the first row, Imagination Policy can learn the multimodal distribution and is equivariant. It realizes around $2^{\degree}\sim 3^{\degree}$ average rotation error and $1\mathrm{cm}$ translation error with different configurations of $P_{a}$ and $P_{b}$; (2) Without downsampling or color information, the rotation error slightly increases; (3) Without data augmentation in training, the performance decreases dramatically since the model cannot learn rotation-invariant features. (4) Compared with the results in the last two rows, the PVCNN-based point cloud encoder outperforms PointNet [41] and the pre-trained equivariant point cloud encoder from NDF [25]. Note that the pre-trained point cloud encoder consumes enormous 3D point clouds from ShapeNet [42] and makes use of Vector Neuron [29] which is guaranteed to output the rotation invariant feature. We hypothesize that the architecture of Vector Neuron [29] and the standard representation limit its expressivity.
(a) banana crown
(b) mug handle
(c) spoon neck
(d) fork handle
(e) banana body
(f) mug body
(g) spoon tail
(h) fork head
Figure 6: Visualization of pick generation and place generation. Top row: multimodal pick-part training labels for different objects. Mid row: single generation for pick . Bottom row: pair generation for place. The generated point cloud is colored in orange. Note the model takes the randomly rotated and downsampled point cloud as input.
Pick/Single Generation Place/Pair Generation
max $\>\boldsymbol{|}\>$ mean $\>\boldsymbol{|}\>$ min rot error (°$\degree$) trans error (cm) rot error (°$\degree$) trans error (cm)
Imagination Policy 0.16 2.02 4.82 0.43 0.94 1.72 0.36 2.34 7.12 0.55 1.05 1.76
w/o downsample 0.47 7.77 39.55 0.43 0.96 1.77 0.29 9.28 28.94 0.65 1.19 1.85
w/o color 0.76 4.41 16.18 0.18 0.87 2.41 0.39 5.05 21.22 0.40 0.97 1.76
w/o augmentation 17.45 125.26 179.27 0.44 1.50 6.53 49.24 130.60 178.59 1.31 14.46 30.80
PointNet Encoder 0.42 3.30 10.09 0.26 0.88 1.56 0.96 4.82 14.31 0.33 0.98 1.74
Pretrained VN Encoder 0.75 5.24 34.06 0.26 0.80 1.56 0.92 6.01 20.68 0.44 1.04 2.13
Table 2: Ablation Results. We report the minimum, mean, and maximum error for single generation and pair generation over 100 runs with randomly rotated and sampled input.

### v2

@@ Abstract
Humans can imagine goal states during planning and perform actions to match those goals. In this work, we propose Imagination Policy, a novel multi-task key-frame policy network for solving high-precision pick and place tasks. Instead of learning actions directly, Imagination Policy generates point clouds to imagine desired states which are then translated to actions using rigid action estimation. This transforms action inference into a local generative task. We leverage pick and place symmetries underlying the tasks in the generation process and achieve extremely high sample efficiency and generalizability to unseen configurations. Finally, we demonstrate state-of-the-art performance across various tasks on the RLbench benchmark compared with several strong baselines and validate our approach on a real robot.
@@ 1 Introduction
Humans can look at a scene and imagine how it would look with the objects in it rearranged. For example, given a flower and a bottle on the table, we can imagine the flower placed in the bottle. Using this mental picture, we can then manipulate the objects to match the imagined scene. However, most robotic policy learning algorithms shortcut this process and map observations directly to actions ($SE(3)$ poses or displacements) [1, 2, 3, 4, 5, 6]. These approaches lose important information about the desired geometry of the scene and are therefore less sample efficient and less precise than they might be.
@@
Inspired by how humans solve tasks, we propose Imagination Policy which takes two point clouds as input and generates a new point cloud combining the inputs into a desirable configuration using a conditional point flow model (see Figure 1). Given the generated point cloud, we use point cloud registration methods to match the observed input point clouds with the “imagined” scene. This gives rigid body transformations which can be used to command a robot arm to manipulate the objects. Imagination Policy consists of two generative processes, each of which uses the above method. As shown in Figure 1a, the pick generator generates the points of the object positioned relative to the gripper point cloud. The place generator generates a pair of objects rearranged together as shown in Figure 1b. Compared to directly generating actions, this adds many degrees of freedom to the generative process which aids optimization and sensitivity to geometric interactions.
Imagination Policy addresses two key challenges in current multitask manipulation policy learning: high precision manipulation and sample efficient learning. Methods like PerAct [1], RVT [2] and Diffuser Actor [3] struggle to learn high precision manipulation policies such as those required to solve the RLBench [7] tasks Plug-Charger and Insert-Knife. Imagination Policy outperforms on these tasks by enabling the model to reason about detailed geometric interaction, such as how different parts of an object’s surface should be displaced, which in turn facilitates precise reasoning about the movement of a tool tip to align with a mating surface. Our method also excels in sample efficiency, the ability to learn good policies with relatively few expert demonstrations. Because Imagination Policy reasons about the desired relative configuration of two objects, it can more easily incorporate symmetries of two object systems, called bi-equivariance, into the model. This significantly improves the sample efficiency of the model. While previous work [8, 9, 10, 11] has also used this kind of bi-equivariant structure to improve sample efficiency, our method is the first to apply the idea outside of the pick-place setting in the more general key-frame multitask setting. While Imagination Policy can still solve pick-place tasks, it can also solve more general manipulation tasks like Plug-Charger, Insert-Knife, and Open-Microwave.
Our contributions in this paper are as follows. 1) We are the first to propose a generative point cloud model to estimate desired rigid body motion in a keyframe manipulation setting. 2) We show how to implement $SE(3)$ bi-equivariant constraints in this setting. 3) We demonstrate that the resulting method, Imagination Policy, achieves state-of-the-art performance on several RLBench tasks against several strong baselines.
@@ 2 Related Work
Point Cloud Generation. Previous works have explored point cloud generation using VAEs [12, 13] and GANs [14, 15]. Recently, score-based denoising models and normalizing flows [16, 17, 18, 19, 20, 21] have demonstrated the power and flexibility to generate high-quality point clouds. For example, Zhou et al. [18] proposed a probabilistic diffusion approach for unconditional point cloud generation. Luo and Hu [19] and Vahdat et al. [20] formulated conditional point cloud diffusion. PSF [22] achieved fast point cloud generation with rectified flow [21]. Ours, however, generates pick and place point clouds conditioned on the observation that can be used to estimate a rigid action to command the robot arm.
@@
Point Clouds in 3D Pick-and-place Manipulation. Point clouds provide a flexible, geometric representation to encode object shapes and poses. In terms of pick-and-place manipulation, Simeonov et al. [23] used $\mathrm{SE}(3)$-invariant point features to encode descriptor fields enabling sample efficient policy learning. Simeonov et al. [4] extended Diffusion Policy [24] to work with point cloud observations and to learn multimodal actions. Pan et al. [25] propose TAX-POSE which is closely related to our method. Pan et al. [25] used two segmented point clouds as input and directly output a new point cloud using a weighted summary of target points together with residual predictions. Concurrently, Eisner et al. [26] adopted TAX-POSE [25] with relative distance inferred by a kernel method. However, it was designed to output the new point cloud directly in one step without penalty for the generated results. Our method, instead, uses generative models to predict the movement of each point iteratively with a velocity model. Moreover, they are limited to single-task training, only work in one-step pick-and-place settings, and thus cannot be applied to complex tasks without predefined prior actions. Recently, Shridhar et al. [1], Goyal et al. [2], and Ke et al. [3] showed impressive multi-task capabilities with transformer-based architectures. However, these methods require hundreds of expert demonstrations and cannot successfully learn high-precision tasks. In contrast, our method leverages bi-equivariant symmetry and amortizes the action prediction across multiple tasks. As a result, it can solve high-precision tasks with few demonstrations.
@@
Symmetry in Robot Learning. Robotic tasks defined in 3D Euclidean space are invariant to translations, rotations, and reflections which redefine the coordinate frame but do not otherwise alter the task. Recent advancements in equivariant modeling, such as those discussed by [27, 28, 29, 30], offer a convenient approach to capturing these symmetries in robotics. Zhu et al. [31] and Huang et al. [32] utilized equivariant models to enforce pick symmetries for grasp learning. Yang et al. [33] proposed an equivariant policy for deformable and articulated object manipulation on top of pre-trained equivariant visual representation. Other works [23, 34, 8, 35, 9, 10, 11, 36, 37, 38, 39] leverage symmetries in pick and place and achieve high sample efficiency. However, they are limited to single-task pick-and-place equivariance. As a result, they cannot be directly applied to the Plug-Charger and Insert-Knife tasks without a pre-place action. Our proposed method, however, can achieve bi-equivariance in the key-frame and multi-task setting. In addition, we employ equivariant action inference using an invariant point cloud generating process, which is different from previous methods.
@@ 3.1 Pair Generation for Place
The generator input is $p^{k}=(x_{t}^{k},f^{k},f_{\ell},t)$. We then optimize $\theta$ with respect to the loss function defined in Equation 2.
Estimating the Action. Given two sets of points $P_{a}\cup P_{b}=(p_{a}^{1},p_{a}^{2},\cdots,p_{a}^{n})\cup(p_{b}^{1},p_{b}^{2},\cdots,p_{b}^{m})$ and their corresponding target positions $P_{ab}=(T_{a}p_{a}^{1},T_{a}p_{a}^{2},\cdots,T_{a}p_{a}^{n})\cup(T_{b}p_{b}^{1},T_{b}p_{b}^{2},\cdots,T_{b}p_{b}^{m})$, we can recover the rigid transformations $T_{a}$ and $T_{b}$ using SVD [40]. However, the output $\hat{P}_{a}\cup\hat{P}_{b}$ of the generator $f_{\mathrm{gen}}$ is not constrained to be given by rigid transforms of the original two point clouds. Each point may move independently by transformation $T^{i}_{\alpha}$ such that $\hat{P}_{a}\cup\hat{P}_{b}=(T^{1}_{a}p_{a}^{1},T^{2}_{a}p_{a}^{2},\cdots,T_{a}^{n}p_{a}^{n})\cup(T^{1}_{b}p_{b}^{1},T^{2}_{b}p_{b}^{2},\cdots,T^{m}_{b}p_{b}^{m})$. We can still use SVD to estimate the best fitting $\hat{T}_{a}$ between $(P_{a},\hat{P}_{a})$ as well as $\hat{T}_{b}$ between $(P_{b},\hat{P}_{b})$. Assuming $P_{b}$ represents the object to be arranged and $P_{a}$ represents the placement, as shown in Figure 2, the pre-place or place action can be calculated as $a_{\mathrm{place}}=(\hat{T}_{a})^{-1}\hat{T}_{b}$.
Figure 2: Architecture of Imagination Policy. (a). Encoding the observed point features as $F_{a}$ and $F_{b}$. (b). Conditional pair generation of the place scene from random Gaussian noise. $x_{t}^{k}$ illustrates the $k$-th noise at time step $t$ with the point feature $f^{k}$ and $f_{\ell}$ is the language feature. (c). Estimating the rigid transformation ($T_{a}$ and $T_{b}$) from the observed point cloud to the generation using correspondence.
Realizing Bi-equivariance. As noted in prior work [8, 9, 10], place actions that transform an object B with respect to another object A are bi-equivariant. That is, independent transformations of object B with $g_{b}\in\mathrm{SE}(3)$ and object A with $g_{a}\in\mathrm{SE}(3)$ result in a change ($a^{\prime}_{\mathrm{place}}=g_{a}a_{\mathrm{place}}g_{b}^{-1}$) to complete the rearrangement at the new configuration. Leveraging bi-equivariant symmetries can generalize learned place knowledge to different configurations and improve sample efficiency.Our placement model is constrained to be bi-equivariant, with invariant generation during training.
@@
Assuming rotation-invariant Gaussian noise $X_{0}$, if the encoded point feature $F_{a}$ and $F_{b}$ are invariant to rotations then $f_{\mathrm{place}}$ is bi-equivariant
@@
With the same generated points $P_{ab}$, the estimated transformation from rotated observation $g_{a}P_{a}={(g_{a}p_{a}^{1},g_{a}p_{a}^{2},\cdots,g_{a}p_{a}^{n})}$ to $\hat{P}_{a}$ is $\hat{T}_{a}g_{a}^{-1}$. Similarly, the estimated transformation from $g_{b}P_{b}$ to $\hat{P}_{b}$ is $\hat{T}_{b}g_{b}^{-1}$. Then, the new place action $a^{\prime}_{\mathrm{place}}$ can be calculated as $a^{\prime}_{\mathrm{place}}=(\hat{T}_{a}g_{a}^{-1})^{-1}\hat{T}_{b}g_{b}^{-1}=g_{a}\hat{T}_{a}^{-1}\hat{T}_{b}g_{b}^{-1}=g_{a}a_{\mathrm{place}}g_{b}^{-1}$, which satisfies bi-equivariance .∎
@@ 3.2 Single Generation for Pick
Assuming rotation-invariant Gaussian noise $X^{P_{a}}_{0}$ , $f_{\mathrm{pick}}$ is equivariant to rotations on the pick target if the encoded point feature $F_{a}$ and $F_{b}$ are rotation-invariant: $f_{\mathrm{pick}}(P_{a},g_{b}\cdot P_{b})=g_{b}f_{\mathrm{pick}}(P_{a},P_{b})$.
Specifically, if there is a rotation $g_{b}$ acting on $P_{b}$, the generated points $\hat{P}_{b}$ are the same as those without rotation. The estimated transformation from $g_{b}P_{b}$ to $\hat{P}_{b}$ is $\hat{T}_{b}g_{b}^{-1}$ and the new pick action can be calculated as $a^{\prime}_{\mathrm{pick}}=(\hat{T}_{b}g_{b}^{-1})^{-1}=g_{b}\hat{T}_{b}^{-1}$, which realizes the desired equivariance property.
@@ 4 Experiments
Model Architecture Details. The generative models $f_{\mathrm{pick}}$ and $f_{\mathrm{place}}$ share the same architecture. Each has two point cloud encoders and a generation network. We select PVCNN [42] as the backbone of our point encoders, which output a 64-dimension feature for each point. We use a pre-trained CLIP-ViT32 [41] model as our language encoder and project the language embedding to a 32-dimension vector with a linear layer. The time step $t$ is encoded as a 32-dimension positional embedding. We also encode a binary mask that indicates if the point belongs to $P_{a}$ or $P_{b}$ as a 32-dimension positional embedding. As a result, the generator input of a point is a 163-dimension vector. We adopt PSF [22] as our generator backbone. Both $f_{\mathrm{pick}}$ and $f_{\mathrm{place}}$ are trained end-to-end with the MSE loss defined in Equation 2 and Equation 5. We use the Adam optimizer with an initial learning rate of $10^{-4}$. Training takes 7 hours to converge with $200$k training steps on a single RTX-4090 graphic card. During inference, we randomly sample $X_{0}$ from a Gaussian distribution and integrate over $v_{\theta}$ with 1000 steps to generate $P_{ab}$ and calculate the action. Generating one batch takes $20$ seconds.
All point clouds $P_{a}$ and $P_{b}$ in our experiments are captured by RGB-D cameras instead of directly sampling from the ground truth mesh. We first center the point cloud and then downsample by selecting at most one point in each cell of a 4mm voxel grid. We further randomly subsample or duplicate to get $2048$ points for $P_{a}$ and $P_{b}$. To get rotation-invariant generation, we apply extensive $\mathrm{SO}(3)$ data augmentation to $P_{a}$ and $P_{b}$ during training, i.e., an $\mathrm{SO}(3)$ rotation is sampled uniformly at each training step. This enforces $f_{\mathrm{gen}}(g_{a}P_{a},g_{b}P_{b})=f_{\mathrm{gen}}(P_{a},P_{b})$, which leads to the desired symmetry properties from Proposition 1. We found the results slightly outperform the equivariant point encoder of Vector Neurons [28], as shown in Table 3. We hypothesize that VN’s expressivity is not as strong as that of PVCNN.
@@
Baselines. Our method is compared against three strong baselines: PerAct  [1] is the state-of-the-art multi-task behavior cloning agent that tokenizes the voxel grids together with a language description of the task and learns a language-conditioned policy with Perceiver Transformer [43]. RVT  [2] projects the 3D observation onto five orthographic images and uses the dense feature map of each image to generate 3D actions. 3D Diffuser Actor [3] is a variation of Diffusion Policy [24] that denoises noisy actions conditioned on point cloud features. Comparison with this baseline tests the importance of point cloud generation since this baseline generates actions directly. RPDiff [4] consumes segmented $P_{a}$ and $P_{b}$ and denoises the relative pose iteratively. To make a fair comparison, we adapt [4] to a multi-task policy. See Appendix 6.4 for our implementation details. NDFs [23] and its variation [34] are not included since they require per-object pretraining.
Settings. All methods are trained as multi-task models. There are four cameras (front, right shoulder, left shoulder, hand) pointing toward the workspace. For our method, we formulate the action sequence as (pick, preplace, place), as shown in Figure 4. Specifically, our method generates the pick action with $f_{\mathrm{pick}}$, and the preplace and place action with $f_{\mathrm{place}}$ simultaneously. We use the ground truth mask to segment $P_{a}$ and $P_{b}$ for RPDiff [4] and our method, as shown in Figure 4a.
@@
Results. We report the results of all methods in Table 1. Several conclusions can be drawn from Table 1: 1) Imagination Policy significantly outperforms all baselines trained with 10 demos on all the tasks except Put-Plate. It can also achieve over $90\%$ success rates in Phone-on-Base and Stack-Wine. 2) For tasks with a high-precision requirement, e.g., Plug-Charger, Insert-Knife and Put-Roll, Imagination Policy has a relatively high success rate while all the baselines fail to learn a good policy. 3) Imagination Policy achieves better sample efficiency and demonstrates few-shot learning performance. With one or five demonstrations, it sometimes outperforms the baselines trained with 10 demonstrations, e.g., Imagination Policy achieves a $97.2\%$ success rate on Stack-Wine trained with 5 demos while the best baseline can only achieve ${32\%}$. We believe this sample efficiency is due to the models equivariance which allows it exploit the symmetry inherent in the generation task. In the end, our method underperforms one baseline in Put-Plate. We hypothesize that the object in this task is symmetric and is hard to encode with distinguishable point features, which might result in wrong correspondences when estimating the rigid transformations. Since many complex manipulation tasks can be decomposed as a sequence of single pick and place, we illustrate that our method can address long-horizon tasks in Appendix 6.2.